OpenStudy (anonymous):

Write each expression in the standard form for a complex number, a + bi. a. [3(cos(27°)) + isin(27°)]^5 b. [2(cos(40°)) + isin(40°)]^6

OpenStudy (anonymous):

\[\large\begin{align*}\bigg[3(\cos27^\circ+i\sin27^\circ)\bigg]^5&=\left(3e^{27^\circ i}\right)^5\\ &=3^5e^{27\cdot5i}\\ &=3^5e^{135^\circ i}\\ &=3^5\bigg[\cos135^\circ +i\sin135^\circ\bigg]\\ &=3^5\left[-\frac{1}{\sqrt2} +\frac{1}{\sqrt2}i\right]\\ &=-\frac{243}{\sqrt2} +\frac{243}{\sqrt2}i \end{align*}\] (b) can be done similarly.